OpenStudy (anonymous):

Write 1.231231231231... as a fraction

OpenStudy (anonymous):

0.1x= 0.123123123... 100x= 123.123... 99.9x= 123 x=1230/999

OpenStudy (amistre64):

1 + 231231231231... /10^(...) = (10^(...) + 231231231231...)/10^(...) :)